OpenStudy (anonymous):

find solution set may use factoring 5x^4=5,000x

OpenStudy (anonymous):

divide both side by x and see what you get. ok?

OpenStudy (anonymous):

divide both sides by 5 first...

OpenStudy (anonymous):

you'll get x^4 = 1000. now use the formula (a^2 - b^2) = (a+b)(a-b)

OpenStudy (anonymous):

still need help?

OpenStudy (anonymous):

x^4 - 1000 (x^2 - 10sqrt10)(x^2 + 10sqrt10) (x^2 - 10sqrt10) can also be factored...